take 100 set of 10 tosses of an unbiased coin. in how many cases do u expect to get 7 heads at least

Answer by stanbon(75887)

take 100 sets of 10 tosses of an unbiased coin. 1
in how many cases do u expect to get 7 heads at least
---------
For each set of 10 tosses you have a binomial problem
with n = 10 and p(heads) = 1/2
----
P(7<= x <= 10) = 1-P(0<= x <= 6) = 1-binomcdf(10,1/2,6) = 0.1718
------------
Since the probabilities for each of the 100 sets are independent
the expected # of sets with at least 7 heads is 0.1718 *100 = 17
---------------
Cheers,
Stan H.
---------------